OpenStudy (anonymous):

For integers a and b, where 4 (less than) a (less than) 6 and 8 (less than) b (less than) 10, what is the greatest possible value of 3/(b-a)? I know the answer but I need an explanation as to how I get the answer. Answer: 3/2

OpenStudy (mathmath333):

can u form the equation from the words

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[4\le a \le 6\] \[8\le b \le 10\]

OpenStudy (mathmath333):

it should be \(\large\tt \begin{align} \color{black}{ 4<a<6\hspace{.33em}\\~\\ 8<b<10\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\Huge \leq \) is pronounced as "less than equal to"

OpenStudy (anonymous):

oh yes thats what i meant my fault

OpenStudy (mathmath333):

ohk

OpenStudy (mathmath333):

i mean \(\Large \dfrac{3}{b-a}_{max}\)

OpenStudy (anonymous):

Alright yes thats what the question is asking for

OpenStudy (anonymous):

The answer its says is 3/2

OpenStudy (anonymous):

But I'm not sure how to get there exactly

OpenStudy (mathmath333):

ohk thats easy

OpenStudy (mathmath333):

for finding \(\Large \dfrac{3}{b-a}_{max}\) do u agree that \(\large (b-a)\) should be minimum

OpenStudy (anonymous):

Hmm how is that so?

OpenStudy (mathmath333):

well consider the fraction \(\dfrac{8}{y}\) if u have y digits from \({1~to~9}\) when will be \(\ \dfrac{8}{y}\) will be maximim ,try to put each digit between 1 and 9

OpenStudy (anonymous):

Okay i think i got you now

OpenStudy (mathmath333):

so which digit of \(y\) will make \(\dfrac{8}{y}\)maximum

OpenStudy (anonymous):

Since it's between 1 and 9 well 2 cause itll give you 4 right?

OpenStudy (mathmath333):

okay if \(y\) was \(1\) then ?

OpenStudy (anonymous):

8 would be the max

OpenStudy (mathmath333):

so for making fraction maximum the denominator should be the least did u agree that ?

OpenStudy (anonymous):

Yes!

OpenStudy (mathmath333):

so or finding \(\Large \dfrac{3}{b−a}_{max}\) \(\Large (b−a)\) should be minimum, right ?

OpenStudy (anonymous):

Yes alright now i understand your poibt.

OpenStudy (mathmath333):

ohk so now for making \(\Large (b-a)\) minimum do u think \(\large b\) should be minimum and \(\large a\) should be maximim or do u think \(\large a\) should be minimum and \(\large b\) should be maximim

OpenStudy (anonymous):

b minimum and a maximum

OpenStudy (mathmath333):

thats right !

OpenStudy (anonymous):

3/8-6 correct?

OpenStudy (anonymous):

3/2

OpenStudy (mathmath333):

yes! \(\Huge \checkmark\)

OpenStudy (anonymous):

Wow so this was a really simple equation hah XD Alright thank you so much!

OpenStudy (mathmath333):

yw